# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 5.1.2 Order parameters
where $V(\bm{x},{\bm{y}})$ is the lattice Coulomb potential associated with a unit charge located at ${\bm{y}}$, and $E_{\mu}({\bm{x}},{\bm{y}})=V({\bm{x}}+\hat{\mu},{\bm{y}})-V({\bm{x}},{\bm{y}})$. Note that this operator cannot be defined for periodic boundary conditions, as $V(\bm{x},{\bm{y}})$ does not exist for these boundary conditions. A rigorous definition in a finite system is possible instead, if C boundary conditions are used. [^21] The lattice Coulomb potential $V({\bm{x}}-{\bm{y}})$ is the solution of the lattice Poisson equation $\sum_{\mu}\Delta^{-}_{\mu}\Delta_{\mu}V({\bm{x}},{\bm{y}})=-\delta_{{\bm{x}},{\bm{y}}}$, where the lattice derivatives act on the ${\bm{x}}$ variable. A consistent definition of $\bm{\Gamma}_{\bm{x}}$ requires the use of boundary conditions for which the lattice Poisson equation has a unique solution. $C^{*}$ boundary conditions, but not periodic boundary conditions, have this property. The operator ${\bm{\Gamma}}_{\bm{x}}$ is invariant under local gauge transformations, but transforms as ${\bm{\Gamma}}_{\bm{x}}\to e^{i\varphi}{\bm{\Gamma}}_{\bm{x}}$ under a global U(1) transformation ${\bm{z}}_{\bm{x}}\to e^{i\varphi}{\bm{z}}_{\bm{x}}$, thus it behaves as a charged gauge-invariant operator. [^22] We are using here $C^{*}$ boundary conditions, so that ${\bm{z}}_{\bm{x}}\to e^{i\varphi}{\bm{z}}_{\bm{x}}$, $A_{{\bm{x}},\mu}\to A_{{\bm{x}},\mu}$ is not a gauge transformation (it would be for periodic boundary conditions). Indeed, if we define $\Lambda_{\bm{x}}=\varphi$, $A_{{\bm{x}},\mu}$ transforms nontrivially on the boundaries due the fact that $\Lambda_{\bm{x}}$ is antiperiodic.
@@
The Lorenz-gauge representation of the charged operator ${\bm{\Gamma}}_{\bm{x}}$ in terms of the local scalar field ${\bm{z}}_{\bm{x}}$ allows us to use the standard RG framework, which applies to local operators [6, 113, 125, 135, 95]. In particular, we can predict the critical behavior of its correlation functions. In particular, we can define a critical exponent $\eta_{\Gamma}$ from the critical-point relation
@@ 5.2.1 The Coulomb and Higgs phases
As shown in the left panel of Fig. 2, the one-component LAH model presents only two phases: a Coulomb (C) phase, in which gauge correlators are gapless, and a Higgs (H) phase in which gauge correlators are gapped; see, e.g., Ref. [50]. The C and H phases can also be characterized by the confinement/deconfinement of charged gauge-invariant excitations [46, 47, 50], as we discuss below. The C and H phases are separated by a transition line connecting the transition points occurring in the $J\to\infty$ and $\kappa\to\infty$ limits, where the noncompact LAH model becomes equivalent to the 3D IXY model discussed in Sec. 3.5.1 and to the standard O(2)-vector spin model, respectively. As discussed in Ref. [93] and reviewed below, the critical behavior along the whole CH line is the same as that of the IXY model.
@@ 5.2.2 IXY critical behavior along the Coulomb-Higgs transition line
In the limit $\kappa\to\infty$, all plaquettes $F_{{\bm{x}},\mu\nu}$ vanish. Thus, in infinite volume we can set $A_{{\bm{x}},\mu}=0$ everywhere up to a gauge transformation, obtaining the XY model. Therefore, the one-component LAH model is expected to undergo an XY transition at [288, 289, 290] $J_{c}(\kappa\to\infty)=0.22708234(9)$ for $\kappa\to\infty$ [this XY transition is denoted by O(2) in Fig. 2]. Note that, as discussed in Sec. 3.5.1, the continuous transitions at the two endpoints have a different nature and belong to two different universality classes, even if related by duality [42, 65]. Indeed, the duality transformations only relate thermal quantities.
The behavior along the CH line has been discussed in Ref. [93], finding that continuous transitions along the CH line belong to the same universality class as that of the IXY model, obtained for $J\to\infty$. Note that it is not possible to have finite-$\kappa$ transitions with the same critical features as those of the spin XY model obtained in the limit $\kappa\to\infty$. Indeed, gauge interactions are a relevant perturbation of the spin XY model. As discussed in Sec. 4.1.2 the uncharged O(2) FP associated with the spin XY universality class is unstable with respect to a perturbation proportional to the gauge coupling [the corresponding eigenvalue of the stability matrix is negative, see Eq. (33)].
The IXY behavior of the transitions is confirmed by the FSS behavior of the energy cumulants [93, 65], see also A.3.4. Results for the third energy cumulant are shown in the left panel of Fig. 3. Data nicely scale if the XY critical exponent $\nu_{\rm XY}=0.6717(1)$ is used. Moreover, the scaling curves match the curve obtained for the IXY model, represented as a solid line in the left panel of Fig. 3. These results confirm that all CH transitions belong to the IXY, or gauge XY, universality class.
Figure 3: (Adapted from Ref. [93]) Scaling plot of the third energy cumulant, defined by $C_{3}=-L^{-3}\langle(H-\langle H\rangle)^{3}\rangle$, for $N=1$ (data across the CH line at fixed $J=1$), $N=2$ (data across the MH line at fixed $J=1$), and $N=25$ ((data across the MH line at fixed $J=0.4$), from left to right. The FSS behavior of $C_{3}$ is discussed in A.3.4, see Eq. (155). We plot $c_{B}C_{3}L^{3-3/\nu}$ versus $c_{X}X$, where $X=(\kappa-\kappa_{c})L^{1/\nu}$, setting $\nu=\nu_{\rm XY}\approx 0.6717$, where $\nu_{\rm XY}$ is the XY value of the critical exponents (see B for a list of estimates of $\nu_{\rm XY}$), and $\kappa_{c}$ is the critical-point value. The nonuniversal constants $c_{X}$ and $c_{B}$ are fixed by requiring the data to match the IXY curve (solid line). See Ref. [93] for details. The nice data collapse confirms that all transitions belong to the IXY universality class.
@@
The RG anomalous dimension $\eta_{A}$ of the gauge field has been determined in Ref.[93], by performing a FSS analysis of the gauge correlation function $C_{\mu\nu}({\bm{x}},{\bm{y}})$ defined in Eq. (55) in the Lorenz gauge. The numerical FSS analyses [93] confirm the exact result $\eta_{A}=1$ that can be proved within the IXY model.  [^23] This exact result follows from the correspondence between the zero-momentum $F_{{\bm{x}},\mu\nu}$ correlation function in the IXY model and the helicity modulus $\Upsilon$ computed in the dual XY model [65], and from the fact that $\Upsilon$ scales as $L^{-1}$ in the XY model [291], see Ref. [93] for details.
Note that the critical behavior of the charged scalar correlations along the $N=1$ CH line differs from that of the scalar correlations in the XY model obtained for $\kappa\to\infty$, see Fig. 2. Indeed, charged scalar correlations are characterized by the critical exponent $\eta_{\Gamma}\approx-0.7$ along the CH line at finite $\kappa$, definitely different from the XY value [95] $\eta_{z}\approx 0.038$ obtained for the limiting model at $\kappa=\infty$. Obviously, in a finite volume the Lorenz-gauge correlations of the field ${\bm{z}}_{\bm{x}}$ along the CH line converge to the scalar correlations of the XY model for $\kappa\to\infty$. However, this result does not imply that their asymptotic infinite-size behavior is the same, since the $\kappa\to\infty$ and the $L\to\infty$ limits do not commute: as discussed in Sec. 4.1.2 the uncharged XY FP, which controls the critical behavior for $\kappa\to\infty$, is unstable with respect to gauge fluctuations, giving rise to a nonanalytic crossover behavior.
@@
We finally remark that, since the RG flow of the AHFT in the $\varepsilon$-expansion does not provide stable 3D charged FPs for $N=1$, see Sec. 4.1.2, the IXY critical behavior along the CH transition line is apparently unrelated with the RG flow of the AHFT. However, as discussed above, the IXY model has an effective description in terms of a model which appears to be equivalent to the AHFT in the Lorenz gauge, with the same field content and global symmetries. To reconcile this interpretation of the IXY transition with the results of Sec. 4.1.2, one may conjecture that, for $N=1$, the 3D AHFT RG flow admits a stable FP that is not smoothly connected with the large-$N$ and $\varepsilon$-expansion regimes: as $N$ or the dimension $d$ increases, the FP apparently disappears, so it cannot be identified in four dimensions or for large $N$. This field-theoretical interpretation is also consistent with the exact result $\eta_{A}=1$ for the IXY universality class, since $\eta_{A}=1$ holds at any charged FP of the 3D AHFT, see Eq. (32). An analogous conjecture has been put forward for other statistical systems, for example, for some frustrated spin systems. In that case 3D field-theoretical high-order perturbative analyses provided evidence of stable FPs not analytically connected with those appearing close to four dimensions and in the large-$N$ limit of the 3D theory [95, 158, 161]. Field-theoretical approaches to describe the 3D critical behavior of $N=1$ systems (superconductors) were also discussed in Refs. [292, 60, 293, 61].
@@ 5.3.1 The Coulomb, Molecular, and Higgs phases
The phase diagram is characterized by three different phases. For small $J$ there is a Coulomb (C) phase, which is SU($N$) symmetric and in which the gauge field is gapless and scalar charged modes are confined. For large $J$ values there are two phases in which the SU$(N)$ symmetry is broken. They are characterized by the different behavior of the gauge and scalar nonlocal charged modes. In the molecular (M) phase the gauge field is long ranged and scalar charges are confined (as in the C phase), while in the Higgs (H) phase gauge fields are gapped and charges deconfined, as it occurs in the one-component model.
@@
The three phases of the multicomponent LAH model are separated by three different transition lines: the CM, CH, and MH transition lines, which start from the transition points located at $\kappa=0$, $J=\infty$ and $\kappa=\infty$, and are expected to meet at one point, as shown in Fig. 2. In particular, the phase diagram in the large-$N$ limit is expected to have a simple shape [93]: The transition lines separating the C, M, and H phases become straight lines: the MH line is given by $\kappa=\kappa_{c}(J=\infty)$, while che CM and CH lines are given by $J=J_{c}(\kappa=0)=J_{c}(\kappa\to\infty)$.
@@
The Higgs phase is a phase where charges deconfine and gauge fields are gapped. Thus, the nonlocal gauge-invariant charged vector operator [286, 91, 93] ${\bm{\Gamma}}_{\bm{x}}$ defined in Eq. (51) is critical both along the CH and the MH line. Analogously, along these two lines also gauge correlation functions (for instance, the correlation function defined in Eq. (55) computed in the Lorenz gauge) are critical.
The transitions along the CM, CH and MH lines may be of first order or continuous and, in the latter case, belong to universality classes that may depend on the number $N$ of scalar components. The continuous transitions are related to the stable (charged or uncharged) FP of the RG flow of the corresponding effective theory, each one with its own attraction domain in the model parameter space.
Figure 5: (Adapted from Ref. [87]). Numerical results for $N=2$ across the CM line. (Left) Plot of $\xi_{Q}/L$, where $\xi_{Q}$ is the second-moment correlation length defined in terms of the correlation function (58) (the FSS behavior of $\xi_{Q}/L$ is discussed in A.3.2), as a function of $(J-J_{c})L^{1/\nu}$ (as a function of $J$ in the inset), with $J_{c}=0.7099$ and $\nu=\nu_{\mathrm{O(3)}}=0.7117$. Data are obtained by varying $J$ for fixed $\kappa=0.04$. The good collapse of the data provides evidence that the transition belongs to the O(3) universality class, as expected. (Right) Plot of the Binder parameter $U_{Q}$ of the operator $Q_{\bm{x}}^{ab}$ [defined in Eq. (57)] as a function of the ratio $\xi_{Q}/L$ (the expected FSS behavior is discussed in A.3.3); we report data obtained by varying $J$ for fixed $\kappa=0.04$ and $\kappa=0$. The excellent collapse of the data for $\kappa=0$ and $\kappa=0.04$ onto the same curve provides a robust evidence that both transitions belong to the same O(3) universality class.
@@ 5.3.2 The Coulomb-Molecular transition line
As discussed in Sec. 3.2.1, for $N\geq 3$ only first-order transitions are possible. Continuous transitions are possible for $N=2$. In this case they belong to the Heisenberg universality class. This is supported by the FSS analyses of numerical results shown in Fig. 5, see A.3 for some details.

### v3

@@ 5.1.2 Order parameters
where $E_{\mu}({\bm{y}},{\bm{x}})=V({\bm{y}}+\hat{\mu},{\bm{x}})-V({\bm{y}},{\bm{x}})$, and $V({\bm{y}},{\bm{x}})$ is the lattice Coulomb potential associated with a unit charge located at ${\bm{x}}$. Note that this operator cannot be defined for periodic boundary conditions, as $V(\bm{y},{\bm{x}})$ does not exist for these boundary conditions. A rigorous definition in a finite system is possible instead, if C boundary conditions are used. [^22] The lattice Coulomb potential $V({\bm{x}},{\bm{y}})$ is the solution of the lattice Poisson equation $\sum_{\mu}\Delta^{-}_{\mu}\Delta_{\mu}V({\bm{x}},{\bm{y}})=-\delta_{{\bm{x}},{\bm{y}}}$, where the lattice derivatives act on the ${\bm{x}}$ variable. A consistent definition of $\bm{\Gamma}_{\bm{x}}$ requires the use of boundary conditions for which the lattice Poisson equation has a unique solution. $C^{*}$ boundary conditions, but not periodic boundary conditions, have this property. The operator ${\bm{\Gamma}}_{\bm{x}}$ is invariant under local gauge transformations, but transforms as ${\bm{\Gamma}}_{\bm{x}}\to e^{i\varphi}{\bm{\Gamma}}_{\bm{x}}$ under a global U(1) transformation ${\bm{z}}_{\bm{x}}\to e^{i\varphi}{\bm{z}}_{\bm{x}}$, thus it behaves as a charged gauge-invariant operator. [^23] We are using here $C^{*}$ boundary conditions, so ${\bm{z}}_{\bm{x}}\to e^{i\varphi}{\bm{z}}_{\bm{x}}$, $A_{{\bm{x}},\mu}\to A_{{\bm{x}},\mu}$ is not a gauge transformation (it would be for periodic boundary conditions). Indeed, if we set $\Lambda_{\bm{x}}=\varphi$, $A_{{\bm{x}},\mu}$ transforms nontrivially on the boundaries due the fact that $\Lambda_{\bm{x}}$ is antiperiodic.
@@
The Lorenz-gauge representation of the charged operator ${\bm{\Gamma}}_{\bm{x}}$ in terms of the local scalar field ${\bm{z}}_{\bm{x}}$ allows us to use the standard RG framework, which applies to local operators [6, 114, 126, 136, 96]. In particular, we can predict the critical behavior of its correlation functions. We define a critical exponent $\eta_{\Gamma}$ from the critical-point relation
@@ 5.2.1 The Coulomb and Higgs phases
As shown in the left panel of Fig. 2, the one-component LAH model presents only two phases: a Coulomb (C) phase, in which gauge correlators are gapless, and a Higgs (H) phase in which gauge correlators are gapped; see, e.g., Ref. [50]. They can also be characterized by the confinement/deconfinement of charged gauge-invariant excitations [46, 47, 50], as we discuss below. The C and H phases are separated by a transition line connecting the transition points occurring in the $J\to\infty$ and $\kappa\to\infty$ limits, where the noncompact LAH model becomes equivalent to the 3D IXY model discussed in Sec. 3.5.1 and to the standard O(2)-vector spin model, respectively. As discussed in Ref. [93] and reviewed below, the critical behavior along the whole CH line is the same as that of the IXY model.
@@ 5.2.2 IXY critical behavior along the Coulomb-Higgs transition line
In the limit $\kappa\to\infty$, all plaquettes $F_{{\bm{x}},\mu\nu}$ vanish. Thus, in infinite volume we can set $A_{{\bm{x}},\mu}=0$ everywhere up to a gauge transformation, obtaining the XY model. Therefore, the one-component LAH model is expected to undergo an XY transition at [163, 154, 151] $J_{c}(\kappa\to\infty)=0.22708234(9)$ for $\kappa\to\infty$ [this XY transition is denoted by O(2) in Fig. 2]. Note that, as discussed in Sec. 3.5.1, the continuous transitions at the two endpoints have a different nature and belong to two different universality classes, even if related by duality [42, 65]. Indeed, duality transformations only relate thermal quantities.
The behavior along the CH line has been discussed in Ref. [93], finding that continuous transitions along the CH line belong to the same universality class as that of the IXY model, obtained for $J\to\infty$. Note that it is not possible to have finite-$\kappa$ transitions with the same critical features as those of the spin XY model obtained in the limit $\kappa\to\infty$. Indeed, gauge interactions are a relevant perturbation of the spin XY model. As discussed in Sec. 4.1.2, the uncharged O(2) FP associated with the spin XY universality class is unstable with respect to a perturbation proportional to the gauge coupling [the corresponding eigenvalue of the stability matrix is negative, see Eq. (33)].
The IXY behavior of the transitions is confirmed by the FSS behavior of the energy cumulants [93, 65], see also A.3.2. Results for the third energy cumulant are shown in the left panel of Fig. 3. Data nicely scale if the XY critical exponent $\nu_{\rm XY}=0.6717(1)$ is used. Moreover, the scaling curves match the curve obtained for the IXY model, reported as a solid line in the left panel of Fig. 3. These results confirm that all CH transitions belong to the IXY, or gauge XY, universality class.
Figure 3: (Adapted from Ref. [93]) Scaling plot of the third energy cumulant, defined by $C_{3}=-L^{-3}\langle(H-\langle H\rangle)^{3}\rangle$, for $N=1$ (data across the CH line at fixed $J=1$), $N=2$ (data across the MH line at fixed $J=1$), and $N=25$ (data across the MH line at fixed $J=0.4$), from left to right. The FSS behavior of $C_{3}$ is discussed in A.3.2, see Eq. (147). We plot $c_{B}C_{3}L^{3-3/\nu}$ versus $c_{X}X$, where $X=(\kappa-\kappa_{c})L^{1/\nu}$, setting $\nu=\nu_{\rm XY}\approx 0.6717$, where $\nu_{\rm XY}$ is the XY value of the critical exponents (see B for a list of estimates of $\nu_{\rm XY}$) and $\kappa_{c}$ is the critical-point value. The nonuniversal constants $c_{X}$ and $c_{B}$ are fixed by requiring the data to match the IXY curve (solid line). See Ref. [93] for details. The nice data collapse confirms that all transitions belong to the IXY universality class.
@@
The RG anomalous dimension $\eta_{A}$ of the gauge field has been determined in Ref.[93], by performing a FSS analysis of the gauge correlation function $C_{\mu\nu}({\bm{x}},{\bm{y}})$ defined in Eq. (58) in the Lorenz gauge. The numerical FSS analyses [93] confirm the exact result $\eta_{A}=1$ that can be proved in the IXY model. [^24] This exact result follows from the correspondence between the zero-momentum $F_{{\bm{x}},\mu\nu}$ correlation function in the IXY model and the helicity modulus $\Upsilon$ computed in the dual XY model [65], and from the fact that $\Upsilon$ scales as $L^{-1}$ in the XY model [319], see Ref. [93] for details.
Note that the critical behavior of the charged scalar correlations along the CH line differs from that of the scalar correlations in the XY model obtained for $\kappa\to\infty$, see Fig. 2. Indeed, charged scalar correlations are characterized by the critical exponent $\eta_{\Gamma}\approx-0.7$ along the CH line at finite $\kappa$, definitely different from the XY value [96] $\eta_{z}\approx 0.038$ obtained for the limiting model at $\kappa=\infty$. Obviously, in a finite volume the Lorenz-gauge correlations of the field ${\bm{z}}_{\bm{x}}$ along the CH line converge to the scalar correlations of the XY model for $\kappa\to\infty$. However, this result does not imply that their asymptotic infinite-size behavior is the same, since the $\kappa\to\infty$ and the $L\to\infty$ limits do not commute: as discussed in Sec. 4.1.2 the uncharged XY FP, which controls the critical behavior for $\kappa\to\infty$, is unstable with respect to gauge fluctuations, giving rise to a nonanalytic crossover behavior.
@@
We finally remark that, since the RG flow of the AHFT in the $\varepsilon$-expansion does not show any evidence of the existence of a stable charged FP for $N=1$, see Sec. 4.1.2, the IXY critical behavior along the CH transition line is apparently unrelated with the RG flow of the AHFT. However, as discussed above, the IXY model has an effective description in terms of a model which appears to be equivalent to the AHFT in the Lorenz gauge, with the same field content and global symmetries. To reconcile this interpretation of the IXY transition with the results of Sec. 4.1.2, one may conjecture that, for $N=1$, the 3D AHFT RG flow admits a stable FP that is not smoothly connected with the large-$N$ and $\varepsilon$-expansion regimes: as $N$ or the dimension $d$ increases, the FP apparently disappears, so it cannot be identified in four dimensions or for large $N$. This field-theoretical interpretation is also consistent with the exact result $\eta_{A}=1$ for the IXY universality class, since $\eta_{A}=1$ holds at any charged FP of the 3D AHFT, see Eq. (32). An analogous conjecture has been put forward for other statistical systems, for example, for some frustrated spin systems. In that case 3D field-theoretical high-order perturbative analyses provided evidence of stable FPs not analytically connected with those appearing close to four dimensions and in the large-$N$ limit of the 3D theory [96, 189, 192]. Field-theoretical approaches appropriate to describe the 3D critical behavior of $N=1$ systems (superconductors) were also discussed in Refs. [320, 54, 60, 321, 61].
@@ 5.3.1 The Coulomb, Molecular, and Higgs phases
The phase diagram is characterized by three different phases. For small $J$ there is a Coulomb (C) phase, which is SU($N$) symmetric and in which the gauge field is gapless and charged scalar modes are confined. For large $J$ values there are two phases in which the SU$(N)$ symmetry is broken. They are characterized by the different behavior of the gauge and nonlocal charged modes. In the molecular (M) phase the gauge field is long ranged and scalar charges are confined (as in the C phase), while in the Higgs (H) phase gauge fields are gapped and charges are deconfined, as it occurs in the one-component model.
@@
The three phases of the multicomponent LAH model are separated by three different transition lines, the CM, CH, and MH transition lines, which start from the transition points located at $\kappa=0$, $J=\infty$, and $\kappa=\infty$, and are expected to meet at one point, as shown in Fig. 2. In particular, the phase diagram in the large-$N$ limit is expected to have a simple shape [93]: The transition lines separating the C, M, and H phases become straight lines: the MH line is given by $\kappa=\kappa_{c}(J=\infty)$, while the CM and CH lines are given by $J=J_{c}(\kappa=0)=J_{c}(\kappa\to\infty)$.
@@
The Higgs phase is a phase in which charges deconfine and gauge fields are gapped. Thus, the nonlocal gauge-invariant charged vector operator [313, 91, 93] ${\bm{\Gamma}}_{\bm{x}}$ defined in Eq. (54) is critical along both the CH and the MH line. Analogously, along these two lines also gauge correlation functions (for instance, the correlation function defined in Eq. (58) computed in the Lorenz gauge) are critical.
The transitions along the CM, CH, and MH lines may be of first order or continuous and, in the latter case, belong to universality classes that may depend on the number $N$ of scalar components. The continuous transitions are related to the stable (charged or uncharged) FP of the RG flow of the corresponding effective theory, each one with its own attraction domain in the model parameter space.
Figure 5: (Adapted from Ref. [87]). Numerical results for $N=2$ across the CM line. (Left) Plot of $\xi_{Q}/L$, where $\xi_{Q}$ is the second-moment correlation length defined in terms of the correlation function (61) (the FSS behavior of $\xi_{Q}/L$ is discussed in A.3.3), as a function of $(J-J_{c})L^{1/\nu}$ (as a function of $J$ in the inset), with $J_{c}=0.7099$ and $\nu=\nu_{\mathrm{O(3)}}=0.7117$. Data are obtained by varying $J$ for fixed $\kappa=0.04$. The good collapse of the data provides evidence that the transition belongs to the O(3) universality class, as expected. (Right) Plot of the Binder parameter $U_{Q}$ of the operator $Q_{\bm{x}}^{ab}$ [defined in Eq. (60)] as a function of the ratio $\xi_{Q}/L$ (the expected FSS behavior is discussed in A.3.4); we report data obtained by varying $J$ for fixed $\kappa=0.04$ and $\kappa=0$. The excellent collapse of the data for $\kappa=0$ and $\kappa=0.04$ onto the same curve provides a robust evidence that both transitions belong to the same O(3) universality class.
@@ 5.3.2 The Coulomb-Molecular transition line
As discussed in Sec. 3.2.1, for $N\geq 3$ only first-order transitions are possible. Continuous transitions are possible for $N=2$. In this case they belong to the Heisenberg universality class. This is supported by the FSS analyses of numerical results shown in Fig. 5.